OpenStudy (anonymous):

simplification of 3^8 over 3^3

OpenStudy (saifoo.khan):

\[\large 3^{8-3}\]

OpenStudy (anonymous):

3^5

OpenStudy (anonymous):

243

OpenStudy (unklerhaukus):

\[3^8 \over 3^3\]\[=3^{8-3}\]\[=3^5\]\[=243\]